During rush hour, Fernando can drive 20 miles using the side roads in the same time that it takes to travel 15 miles on the freeway. If Fernando's rate on the side roads is 9 mi/h faster than his rate on the freeway, find his rate on the side roads.
Answer by checkley71(8405)

20/(X+9)=15/X CROSS MULTIPLY
15X+135=20X
20X-15X=135
5X=135
X=135/5
X=27 IS THE MPH ON THE SIDE ROAD.